Joe has a collection of nickels and dimes that is worth $3.50. if the number of dimes were doubled and the number of nickels were increased by 17, the value of the coin would be $6.45. how many dimes does he have?
Answer by stanbon(75887)

Joe has a collection of nickels and dimes that is worth $3.50.
if the number of dimes were doubled and the number of nickels were increased by 17, the value of the coin would be $6.45.
how many dimes does he have?
-----
Equations:
5n + 10d = 350 cents
5(n+17)+10(2d) = 645 cents
-------------------------------
Simplify:
5n + 10d = 350
5n + 20d = 645-85
---
Subtract 1st from 2nd and solve for "d":
10d = 210
d = 21 (# of dimes)
---
Substitute into 5n+10d = 350
5n + 10(21) = 350
5n + 210 = 350
5n = 140
n = 28 (number of nickels)
============
Cheers,
Stan H.